OpenStudy (anonymous):

The y-intercept of the line whose equation is x - y = -1 is -1 0 (1) correct?

OpenStudy (anonymous):

yes that is correct, the y-intercept is where x=0 plugging in o for x gives you 0-y=-1 -y=-1 y=1

OpenStudy (anonymous):

thanks. can you help me with 1 more?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

The slope of the line passing through the points (2, 7) and (-4, 8) is -6 -1/2 correct? -1/6

OpenStudy (anonymous):

yes that is correct; m=(y2-y1)/(x2-x1) m=(8-7)/(-4-2) m=-1/6

OpenStudy (anonymous):

you say its -1/6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but you said i was correct

OpenStudy (anonymous):

oh sorry i read it wrong, you were not correct try again using the formula m=(y2-y1)/(x2-x1)